How much 30% acid solution should be added to 10% acid solution to make 100 ml of a 12% acid solution?
Answer by stanbon(75887)

How much 30% acid solution should be added to 10% acid solution to make 100 ml of a 12% acid solution?
--------
Equation:
acid + acid = acid
0.30x + 0.10(100-x) = 0.12*100
-----
30x + 10*100 - 10x = 12*100
20x = 2*100
x = 10 ml (amt of 30% solution needed)
100-x = 90 ml (amt of 10% solution needed)
==============
Cheers,
Stan H.